# TRACT: Regression-Aware Fine-tuning Meets Chain-of-Thought Reasoning for LLM-as-a-Judge

### v1

@@ 6.1 On the Distribution Shift between Annotation vs Self-Generated CoTs
To quantify how different $s_{i}^{*}$ and $\hat{s}_{i}$ are, we compare the root mean square error (RMSE) across two pairs of scores: 1) the predicted scores $\hat{y}_{i}(s_{i}^{*})$ and the ground truth scores $y_{i}^{*}$, and 2) the predicted scores $\hat{y}_{i}(\hat{s}_{i})$ and the ground truth scores $y_{i}^{*}$. Notably, both $\hat{y}_{i}(\hat{s}_{i})$ and $\hat{y}_{i}(s_{i}^{*})$ are generated by the same LLM $p(\cdot)$ conditioned on the same input $x_{i}$ from the training set; the only distinction is the CoT: $\hat{y}_{i}(s_{i}^{*})$ conditioned on $s_{i}^{*}$, and $\hat{y}_{i}(\hat{s}_{i})$ conditioned on $\hat{s}_{i}$.
@@ 6.2 Sensitivity to $\lambda$
TRACT relies on the CoT-RAFT objective in Equation 4, which combines the CE loss over CoTs and the RAFT loss over score predictions. An alternative training strategy could involve a sequential, two-stage fine-tuning process: first, fine-tuning with the CE loss on both CoTs and scores (corresponding to baseline 2); and next, fine-tuning on the same dataset using solely the RAFT objective. In the second stage, while the model’s output still includes CoTs, the CE loss is not computed for them, i.e., the second term in Equation 4 is not used.
We empirically find that this sequential fine-tuning approach significantly degrades the model’s ability to generate coherent and complete CoTs. The resulting model struggles to terminate CoT outputs correctly and sometimes yields very long CoTs. Furthermore, the model frequently fails to generate the string ’ So the overall score is’, which is used to determine where to apply the RAIL predictor for the final score prediction. We show some example CoTs in Table 9 in Appendix C.
@@ 6.4 Scaling the Number of Sampled CoTs
We now analyze the effect of sampling multiple CoTs for TRACT during inference; this is done by sampling multiple $\hat{s}$, calculating $\hat{y}_{\rm CR}$ for each CoT (Eq. 3), and averaging the $\hat{y}_{\rm CR}$. For comparison, we also sample multiple CoTs and average the scores for the following two baselines: the model trained using the CE loss (baseline 2, row B.2 in Table 2), and Prometheus-2-7B. To gauge the significance of the differences across methods, we repeat the experiment by varying the random seed. For each random seed, we sample $K$ CoTs to calculate the score, average the score across four testing dataset, and repeat the above process with four random seeds.
@@
Compared to Prometheus, TRACT shows an advantage under the setting where fewer CoTs are sampled. We note that Prometheus is trained over additional data corresponding to a ranking task, and thus comparing TRACT and Prometheus is not strictly fair, disadvantaging TRACT. For this reason, we find it a very strong outcome that TRACT outperforms Prometheus under the limited CoT sampling setting.
@@ 7.1 LLM-as-a-Judge
LLM-as-a-judge paradigm is an application of LLMs as an automatic judge to assess texts (Chiang and Lee, 2023a). There are two types of LLM-as-a-judge: (1) Direct assessment: the LLM assigns a score to a given sample (Chiang and Lee, 2023a; Liu et al., 2023). (2) Pairwise ranking: the LLM determines which sample is better given a pair of samples (Zheng et al., 2023; Wang et al., 2024; Li et al., 2024). In this paper, we focus on improving the direct assessment ability of the LLM.
While strong closed-source LLMs, including GPT-3.5/4 (OpenAI, 2023) and Claude (Anthropic, 2024), show promising evaluation ability, their black-box nature makes them somewhat problematic for open research. Many prior works focus on fine-tuning open-source LLMs into better LLM judges using open-source data (Vu et al., 2024; Li et al., 2024). LLM-as-a-judge is often trained with the CE loss over CoT reasoning and a score (Kim et al., 2024a, b).
@@ 7.2 Training LLMs with Self-Generated Data
Most related to our work is Yang et al. (2024), who argue that self-generated data helps fine-tuning by reducing the distribution gap between the training data and the seed LLM. Ren et al. (2024) show that self-generated data yields lower perplexity under the seed LLM, which can explain the performance gain of fine-tuning on self-generated data. While we also find that self-generated data has a lower perplexity under the seed LLM (Appendix B.4), we use a different fine-tuning objective from prior works. Interestingly, we find that when fine-tuning over self-generated data with the CE loss, as done in prior works, the performance does not improve. On the other hand, our proposed objective benefits from self-generated CoTs significantly, showing how self-generated data is particularly important for TRACT.
@@ 8 Conclusion
In this paper, we introduce Two-stage Regression-Aware fine-tuning with CoT reasoning (TRACT), a method integrating CoT reasoning with the RAFT framework to harness the step-by-step reasoning ability of LLMs while learning the numerical structure of the scoring task. TRACT consists of (1) CoT-RAFT for training, (2) CoT-RAIL for inference, and (3) two-stage fine-tuning for self-generated CoTs training. Our experiments across two models and four datasets show that TRACT consistently outperforms all prior baselines finetuned on the same dataset by a significant margin. Under limited inference compute, TRACT also outperforms Prometheus-2-7B, a same-sized SoTA model trained on more data. We carry out careful ablations to explicate the importance of each component in TRACT. We believe our proposed method and open-source models can benefit the community working on LLM-as-a-judge. In future work, it would be interesting to combine TRACT with pairwise ranking tasks Kim et al. (2024a).
@@ B.1 Quality of the Self-Generated CoTs
###Task Description:
An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing an evaluation criterion is given. (...)
@@
[user_instruction]
@@
[response_to_evaluate]
@@
[reference_answer]
###Score Rubrics: [rubrics]
@@
(a) Input prompt template
The response offers practical, original, and effective solutions (...). So the overall score is 5
@@
Table 4: The input prompt template and an example output used in our main experiment.
@@ B.4 Perpexity of Training Data
Id CoT Train Data Inf $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
Baselines
B.1 CE GPT-4 Decode 0.857 0.857 0.433 0.435 0.423 0.400 0.556 0.541 0.567 0.558
B.2 CE GPT-4 Decode 0.834 0.835 0.484 0.475 0.483 0.467 0.494 0.466 0.574 0.561
B.3 RAIL 0.689 0.683 0.445 0.412 0.487 0.485 0.583 0.547 0.551 0.532
B.4 RAFT GPT-4 RAIL 0.918 0.920 0.493 0.506 0.541 0.509 0.614 0.618 0.642 0.639
TRACT (ours)
6 C-RAFT Self C-RAIL 0.917 0.920 0.493 0.500 0.650 0.605 0.639 0.672 0.675 0.674

### v2

@@ 6.1 On the Distribution Shift between Annotation vs Self-Generated CoTs
To quantify how different $s_{i}^{*}$ and $\hat{s}_{i}$ are, we compare the root mean square error (RMSE) across two pairs of scores: 1) the predicted scores $\hat{y}_{i}(s_{i}^{*})$ and the ground truth scores $y_{i}^{*}$, and 2) the predicted scores $\hat{y}_{i}(\hat{s}_{i})$ and the ground truth scores $y_{i}^{*}$. Notably, both $\hat{y}_{i}(\hat{s}_{i})$ and $\hat{y}_{i}(s_{i}^{*})$ are generated by the same LLM $p(\cdot)$ conditioned on the same input $x_{i}$ in the training set; the only distinction is the CoT: $\hat{y}_{i}(s_{i}^{*})$ conditioned on $s_{i}^{*}$, and $\hat{y}_{i}(\hat{s}_{i})$ conditioned on $\hat{s}_{i}$.
@@ 6.2 Sensitivity to $\lambda$
TRACT relies on the CoT-RAFT objective from Eq. 4, combining the CE loss over CoTs and the RAFT loss over score predictions. An alternative training strategy could involve a sequential, two-stage fine-tuning process: first, fine-tuning with the CE loss over both CoTs and scores (corresponding to baseline 2); and next, fine-tuning on the same dataset using solely the RAFT objective over scores. In the second stage, while the model’s output still includes CoTs, the CE loss is not computed for them, i.e., the second term in Eq. 4 is not used.
We empirically find that this sequential fine-tuning approach significantly degrades the model’s ability to generate coherent and complete CoTs. The resulting model struggles to terminate CoT outputs correctly and sometimes yields very long CoTs. We show some example CoTs in Table 9 in Appendix C. This deficiency highlights the critical role of multi-objective fine-tuning: jointly optimizing both the CE loss for CoTs and the squared-error loss for scores.
@@ 6.4 Scaling the Number of Sampled CoTs
We analyze the effect of sampling multiple CoTs for TRACT during inference; this is done by sampling multiple $\hat{s}$, calculating $\hat{y}_{\rm CR}$ for each CoT (Eq. 3), and averaging the $\hat{y}_{\rm CR}$. For comparison, we also sample multiple CoTs and average the scores for the following two baselines: the model trained using the CE loss (baseline 2, row B.2 in Table 2), and Prometheus-2-7B. To gauge the significance of the differences across methods, we repeat the experiment by varying the random seed. For each seed, we sample $K$ CoTs to calculate the score, average the score across four testing dataset, and repeat the above process with four seeds.
@@
TRACT outperforms Prometheus when sampling fewer CoTs. Given that Prometheus is trained with more data, we find it to be a very strong outcome that TRACT outperforms Prometheus under the limited CoT sampling setting.
7 Comparing Point-Wise LLM-as-a-Judge and Reward Models
Given an instruction and a response, a reward model (RM) assigns a scalar reward to indicate the quality of the response (Wang et al., 2024a; Liu et al., 2024; Dorka, 2024). While this formulation is similar to pointwise LLM-as-a-judge, RMs are typically trained to evaluate specific attributes, e.g., helpfulness, harmlessness, etc. (Bai et al., 2022). As a result, RMs cannot take fine-grained evaluation metrics and evaluate based on the evaluation rubrics. Furthermore, the reward of a RM is only meaningful when comparing the reward of two responses for the same instruction, since RMs are trained with pairwise data. In contrast, the score of point-wise LLM-as-a-judge is meaningful standalone since it directly corresponds to the description of a specific score in the scoring rubrics.
Critique-out-loud (CLoud(Ankner et al., 2024) RM is an RM that generates some CoT reasoning before predicting the reward using a regression head. While CLoud uses self-generated CoT reasoning to improve the numeric prediction (the reward), there are several differences between CLoud and TRACT. First, CLoud is trained on pairwise reward modeling datasets, while TRACT is trained on a point-wise LLM-as-a-judge dataset. Next, TRACT uses the language model head to predict the score, while CLoud uses a regression head to predict the score. Most importantly, CLoud, as a reward model, cannot take fine-grained evaluation rubrics and evaluate the responses based on the evaluation criteria, which is different from all other LLM-as-a-Judge based baselines in Section 4, which can take fine-grained evaluation rubrics.
7.1 Results on Point-wise LLM-as-a-Judge Datasets
Since CLoud also uses CoT to improve numeric predictions, we compare CLoud with TRACT on point-wise LLM-as-a-judge datasets to understand whether RMs can perform well on point-wise LLM-as-a-judge datasets. In row 7 in Table 2, we show the results of CLoud on LLM-as-a-judge datasets. We can see that CLoud performs poorly, and TRACT significantly outperforms CLoud. This is not surprising, as CLoud is a reward model trained on pairwise ranking datasets, and it cannot take the fine-grained evaluation metrics. Our goal here is not to criticize the performance of CLoud; we use the above evidence to highlight that RMs are not the same as point-wise LLM-as-a-judge models.
7.2 Results on RewardBench
Having seen that RM cannot yield reasonable performance on point-wise LLM-as-a-judge datasets, one may ask whether the reverse holds as well: Can point-wise LLM-as-a-judge models be used as a reward model? To answer this question, we use RewardBench (Lambert et al., 2024), a reward model dataset, to evaluate our baseline models and TRACT. Each evaluating instance in RewardBench consists of an instruction, a winning response, and a losing one; the goal is to see whether a reward model can assign a higher reward to the winning response. To test point-wise LLM-as-a-judge models on RewardBench, we prompt the model with both responses separately, obtaining scores for them, and then compare the scores for the final ranking. We report the accuracy on RewardBench.
The results for RewardBench are presented in Table 10 in the Appendix. We see that TRACT performs quite well: the model fine-tuned from Mistral reaches an average score of 0.736, slightly worse than CLoud (0.759). This shows that even though TRACT, a point-wise LLM-as-a-judge model, is not trained on pairwise ranking for reward modeling, it can still obtain reasonably good performance on RewardBench, a pairwise ranking benchmark. This is in stark contrast with the result of CLoud on point-wise LLM-as-a-judge shown in Table 2, where CLoud underperforms TRACT by a large margin. Lastly, we note that TRACT, a model trained on only a point-wise dataset, outperforms Prometheus-2-7B, a model trained on point-wise scoring and pairwise ranking datasets.
@@ 8.1 LLM-as-a-Judge
LLM-as-a-judge paradigm is an application of LLMs as an automatic judge to assess texts (Chiang and Lee, 2023a). There are two types of LLM-as-a-judge: (1) Direct assessment: the LLM assigns a score to a given sample (Chiang and Lee, 2023a; Liu et al., 2023). (2) Pairwise ranking: the LLM determines which sample is better given a pair of samples (Zheng et al., 2023; Wang et al., 2024b; Li et al., 2024). In this paper, we focus on improving the direct assessment ability of the LLM.
Many prior works focus on fine-tuning open-source LLMs into better LLM judges using open-source data (Vu et al., 2024; Li et al., 2024). LLM-as-a-judge is often trained with the CE loss over CoT reasoning and a score (Kim et al., 2024a, b).
@@ 8.2 Training LLMs with Self-Generated Data
Most related to our work is Yang et al. (2024), who argue that self-generated data helps fine-tuning by reducing the distribution gap between the training data and the seed LLM. Ren et al. (2024) show that self-generated data yields lower perplexity under the seed LLM, which can explain the performance gain of fine-tuning on self-generated data. While we confirm that self-generated data has a lower perplexity under the seed LLM (Appendix B.4), we use a fine-tuning objective different from prior works. Interestingly, we find that fine-tuning over self-generated data with the CE loss, as recommended in prior works, does not improve the performance. Our proposed objective does significantly benefit from self-generated CoTs, showing how self-generated data is important for TRACT.
@@ 9 Conclusion
In this paper, we introduce Two-stage Regression-Aware fine-tuning with CoT reasoning (TRACT), a method integrating CoT reasoning with the RAFT framework to harness the step-by-step reasoning ability of LLMs while learning the numerical structure of the scoring task. TRACT consists of (1) CoT-RAFT for training, (2) CoT-RAIL for inference, and (3) two-stage fine-tuning for self-generated CoTs training. Our experiments across two models and five datasets show that TRACT consistently outperforms all prior baselines finetuned on the same dataset by a significant margin. Under limited inference compute, TRACT also outperforms Prometheus-2-7B, a same-sized SoTA model trained on more data. We carry out careful ablations to explicate the importance of each component in TRACT. We believe our proposed method and open-source models can benefit the community working on LLM-as-a-judge.
Limitations
One of the limitations of TRACT is that it requires access to the model’s output probabilities, which is not always possible for proprietary models. Our technique can be readily applied to open-source models.
Acknowledgments
We thank the reviewers for their valuable and constructive feedback. We compare with CLoud and include the results of RewardBench based on the reviewers’ suggestions. Cheng-Han Chiang is supported by a Google PhD Fellowship and a Ph.D. scholarship program by Delta Electronics.
@@ B.1 Quality of the Self-Generated CoTs
An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given.
1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general.
2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric.
3. The output format should look as follows: "Feedback: (write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"
4. Please do not generate any other opening, closing, and explanations.
@@
You are a well-known psychiatrist who has a reputation for being empathetic and understanding. A client comes to you saying they’ve had a really hard day at work. They describe their boss as being overly critical and not listening to their ideas, which has left them feeling frustrated and undervalued. They also mention feeling overwhelmed with the workload and a growing sense of loneliness as they are new to the city. How do you respond to them to make them feel heard and understood, as well as offering helpful advice?
@@
It’s indeed challenging to deal with a difficult boss and to carry a large workload, especially when you are new to a city and haven’t established a supportive network. I would suggest taking a step back to think about your situation and perhaps speak to your boss regarding the issues you’re facing. On the other hand, dealing with a large workload can be managed by prioritizing tasks, and maybe you can discuss it with your superiors too.
In regards to feeling lonely, you might want to explore activities or groups that align with your interests to meet new people. And also, remember to take time for yourself to relax and destress. I’m here to assist you as we navigate through these issues.
@@
I hear that you’re dealing with quite a few challenging situations at once. It must be tough to have your ideas dismissed by your boss and to feel undervalued in your work environment. It’s understandable that such experiences can lead to frustration and a feeling of being overwhelmed.
It’s important to acknowledge these feelings and not to dismiss them. Feeling heard and validated in our experiences can often be the first step towards resolving the challenges we face. So, I want you to know that your feelings are completely valid and it’s okay to be frustrated and overwhelmed.
In terms of dealing with your boss, have you considered seeking a private conversation with them to express your feelings? It’s often beneficial to communicate what you’re experiencing, using "I" statements to express how you feel when your ideas are not taken into consideration. This can lead to mutual understanding and possibly a change in behavior.
About the workload, it might help to prioritize tasks and potentially delegate, if possible. Also, consider discussing your workload with your superiors. There might be a misunderstanding about what’s manageable or they might have suggestions about how to handle the situation.
On the personal front, feeling lonely, especially when you’re new to a city, can be really hard. Seek out opportunities to meet new people, perhaps through hobbies, community activities, or online groups. It might take a bit of time, but gradually, you can build a network of friends and acquaintances.
Remember, it’s perfectly okay to have bad days and it’s important to take care of your mental health. Consider incorporating activities into your daily routine that make you happy and help you unwind. This could be anything from reading, yoga, going for a walk, or even listening to your favorite music.
Please know that you’re not alone in this. I’m here to support you through this challenging time and together, we can work towards resolving these issues.
###Score Rubrics:
[Is the model able to identify and react correctly to the emotional context of the user’s input?]
Score 1: The model utterly fails to grasp the user’s emotional context and responds in an unfitting manner.
Score 2: The model sporadically identifies the emotional context but frequently replies in a manner that doesn’t match the user’s emotional status.
Score 3: The model typically identifies the emotional context and reacts suitably, but occasionally misreads or misjudges the user’s feelings.
Score 4: The model often identifies the emotional context and reacts suitably, with minor cases of misreading or misjudging.
Score 5: The model flawlessly identifies the emotional context of the user’s input and consistently responds in a considerate and empathetic manner.
@@
(a) Input example with the input prompt template.
The response offers practical, original, and effective solutions (...). So the overall score is 3
@@
Table 4: The input prompt template (with an input example) and an example output used in our main experiment.
@@ B.4 Perpexity of Training Data
Train/Inference Configuration FB Bench FLASK Vic. Bench MT Bench Average Id CoT Train Data Inf $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ Baselines B.1 CE GPT-4 Decode 0.857 0.857 0.433 0.435 0.423 0.400 0.556 0.541 0.567 0.558 B.2 CE GPT-4 Decode 0.834 0.835 0.484 0.475 0.483 0.467 0.494 0.466 0.574 0.561 B.3 RAIL 0.689 0.683 0.445 0.412 0.487 0.485 0.583 0.547 0.551 0.532 B.4 RAFT GPT-4 RAIL 0.918 0.920 0.493 0.506 0.541 0.509 0.614 0.618 0.642 0.639 TRACT (ours) 7 C-RAFT Self C-RAIL 0.917 0.920 0.493 0.500 0.650 0.605 0.639 0.672 0.675 0.674
@@ Appendix D License
Id Chat Chat-Hard Safety Reasoning Average B.1 0.592 0.290 0.700 0.355 0.484 B.2 0.629 0.305 0.120 0.368 0.355 B.3 0.595 0.358 0.651 0.312 0.479 B.4 0.883 0.581 0.807 0.828 0.775 B.5.1 0.679 0.423 0.673 0.378 0.538 B.5.2 0.855 0.491 0.771 0.765 0.720 B.6 0.965 0.455 0.754 0.862 0.759 TRACT (ours) 7.M 0.927 0.542 0.759 0.716 0.736 7.L 0.922 0.434 0.799 0.837 0.748 A.1 0.777 0.513 0.569 0.482 0.585 A.2 0.886 0.564 0.773 0.645 0.717 A.3 0.564 0.292 0.385 0.194 0.359 A.4 0.824 0.469 0.700 0.568 0.640
Table 10: Accuracy on RewardBench. The indices are identical to those in Table 2. 7.M and 7.L denote TRACT trained from Mistral and Llama-3-8B, respectively. All models are trained from Mistral, except for B.6 (CLoud) and 7.L, which are based on Llama-3-8B. For Prometheus-2-7B, we report two results since it can be used as a pointwise (B.5.1) and a pairwise (B.5.2) LLM-as-a-judge.